OpenStudy (anonymous):

Find the slope of the line whose equation is 2x 4y = 10. I had this question on one of my math quizzes and i got it wrong. i guessed -1/2 but the answer was just 1/2. what did i do wrong?

OpenStudy (anonymous):

neglected to put a sign between 2x and 4y

OpenStudy (anonymous):

- sorry

OpenStudy (anonymous):

1/2

OpenStudy (anonymous):

2x-4y=10

OpenStudy (anonymous):

Always get it in the form y = mx + b and it's foolproof method.

OpenStudy (anonymous):

To find the slope from an equation, the easiest way is to change it into slope-intercept form: y=mx+b \[2x-4y=10\]\[-4y = 10-2x\]\[y = \frac{1}{2}x-\frac{5}{2}\]The coefficient of x is 1/2, so that's your slope.

OpenStudy (anonymous):

when you divided -2x by -4 you should have gotten 1/2x and 1/2 is the slope your signs must have been off